OpenStudy (anonymous):

Whats the difference between a general solution and an implicit solution of a differential equation.

OpenStudy (amistre64):

general solutions are vague and imprecise; works for a family of stuff expplicit means that we can define one variable in terms of the others implicit means the equation cant really be broken down into something useable

OpenStudy (amistre64):

y^2 - y = sin(x) solve for y is difficult so we say its implicit; or implied that y is a function of x

OpenStudy (amistre64):

y = tan(x)-3x is explicitly stating y as a function of x

OpenStudy (amistre64):

.... of course it helps of i can read ...

OpenStudy (anonymous):

so after you have simplified as much as you can and cant go any further leave the equation as an implicit answer

OpenStudy (amistre64):

ok, a general solution is one of a family; involves arbitrary constants

OpenStudy (amistre64):

correct